OpenStudy (anonymous):

Suppose that the x-intercepts of the graph of y = f(x) are -5 and 3 a) What are the x-intercepts of the graph of y = f(x+2)?

OpenStudy (lgbasallote):

if the x intercepts of f(x) is -5 and 3 that means the factors are (x+5) and (x-3) that means the equation is y = x^2 + 2x - 15 does that help?

OpenStudy (anonymous):

Ok I understand what you did there. But how is that supposed to help me find the x-intercepts of the other equation?

OpenStudy (lgbasallote):

you get f(x+2) you do this by substituting x+2 to every x in what i wrote. for example f(x+2) = (x+2)^2 + 2(x+2) -15 so now simplify that and tell me what you get